OpenStudy (anonymous):

Evaluate the expression 5(1251/3). A. 5 B. 25 C. 125 D. 625

OpenStudy (shamil98):

\[\large 5 \times \frac{ 1251 }{ 3 }\] this?

OpenStudy (anonymous):

is (125^ 1/3)

OpenStudy (shamil98):

oh.

OpenStudy (shamil98):

\[\large 5 \times125^{1/3}\] \[\large 5 \times \sqrt[3]{125}\] \[\large = 5 \times 5\] = 25

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

how did u get 5 on this \[3\sqrt[?]{?}125\]